OpenStudy (anonymous):

Let f(x) = 8x3 – 22x2 – 4 and g(x) = 4x – 3.

OpenStudy (anonymous):

Find f of x over g of x

OpenStudy (cggurumanjunath):

f(x)/g(x)=(8x^3-22x^2-4)/(4x-3)

OpenStudy (cggurumanjunath):

is it clear ?

OpenStudy (anonymous):

@cggurumanjunath Let f(x) = 8x3 – 22x2 – 4 and g(x) = 4x – 3. Find f(x)/g(x) so im getting ' 2x2 – 4x – 3 – 13/4x-3

OpenStudy (cggurumanjunath):

no

OpenStudy (cggurumanjunath):

f(x)/g(x)=(8x^3-22x^2-4)/(4x-3) is the final answer

OpenStudy (cggurumanjunath):

see we can't factorise (8x^3-22x^2-4) further so.

OpenStudy (anonymous):

@CGGURUMANJUNATH 2x2 – 4x – 3 – 13 over quantity 4x minus 3 2x2 – 4x – 3 – quantity of 4 x minus 3 over 13 2x2 – 7x – 1 2x2 – 7x – 5 + quantity of x minus 4 over quantity of 4 x minus 3